OpenStudy (hihi):

What is the exact value of the expression √18 +√48 -√50? Simplify completely

OpenStudy (anonymous):

We need to simplify the radicals. I will do the first two: sqrt(18) = sqrt(9 *2) = 3sqrt(2) sqrt(48) = sqrt(16 *3) = 4sqrt(3) Can you try the third one?

OpenStudy (hihi):

sqrt(50)=sqrt(25x5)=5sqrt(5)

OpenStudy (anonymous):

Also written as \[\sqrt{18} = 3\sqrt{2}\] \[\sqrt{48} = 4\sqrt{3}\]

OpenStudy (anonymous):

Nice! Now we have 3 sqrt (2) + 4sqrt(3) - 5 sqrt(2)

OpenStudy (anonymous):

We can combine the first and third term to get -2 sqrt (2) + 4sqrt(3) for our final answer

OpenStudy (hihi):

Oh, I see now, thank you so much meverett04 :)

OpenStudy (anonymous):

You are welcome